A person inside water of refractive index 4/3 sees the setting sun at about
1.42° to the horizontal
2.48°to the horizontal

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
There is no question part posted. Please check and repost the complete question. You can also attach an image. I will be happy to help you.
Khimraj
4 years ago
 

Refractive index of slab, μ=1.5

Total apparent length in the glass =0.06+0.04=0.10m

Thickness of the slab, t=μ× Apparent length in the glass =1.5×0.10=0.15m